Briefly explain how a firm should set a two two-part tariff when faced with two groups of consumers with very different preferences for the good the firm produces.
Answer #1

Two part tariff:

It can be written as T(q)= a+pq, where a is the fixed fee chand p is the marginal price to be paid.The firm chooses a and p to maximise profits given the demand.

Here p=MC, and then a is set to extract maximum consumer surplus from one of the two buyers. This means a should be equal to the surplus enjoyed by least eager buyer, such that he is indifferent between buying or not, and others will experience net gains from the purchase.

Here is an example :

Let there be two different demanders:

q1= 24-p1

q2=24-2p2

MC=6

Here p1=p2=MC=6, q1=18, q2=12.

Now to find a, demander 2 is the least eager buyer( as when q=0, he is willing to pay only 12 while the other can pay 24)

consumer surplus of 2, = 0.5(12-6)*12 =36

That is the maximal entry fee that might be charged without causing this person to leave the market. Hence, the two part tariff becomes:

T(q)= 36+6q